# Structural Alignment Improves Graph Test-Time Adaptation

### v1

@@ 3 Test Error Analysis
Figure 3: (a) Comparison of neighborhood alignment with $\bm{\gamma}$ from model prediction and Oracle on the CSBM graphs (Deshpande et al., 2018). The top (or bottom) subfigures represents the assignment under neighbor shift (or neighbor shift plus label shift, respectively). Nodes are grouped by the entropy of their soft pseudo labels and the y axis shows the accuracy after assigning $\bm{\gamma}$. Ideally, a correct assignment (red) would lead to near-perfect accuracy. However, the assignment based on pseudo labels is far from optimal (blue). From (b) to (c), the figure with the t-SNE visualization of node representations indicates a model trained from the source domain (CSBM with a label distribution $[0.1,0.3,0.6]$) to the target domain (CSBM with a label distribution $[0.3,0.3,0.3]$). The color of the nodes represents the ground-truth labels. The top subfigures of (b) and (c) show the output given by the GNN encoder while the bottom subfigures show the classifier decision boundaries. (d) Analysis of SNR adjustment $\alpha$ with respect to different layers and node degrees on MAG. Detail discussed in Sec. 5.3.
3.1 Distribution Shifts on Graphs
Distribution shifts on graphs were formally studied in previous GDA works (Wu et al., 2020; Liao et al., 2021; Zhu et al., 2021; Wu et al., 2022; You et al., 2023; Zhu et al., 2023). Following their definition, we categorize shifts in graphs into two types: feature shift and structure shift. For simplicity, our analysis is based on a data generation process: ${\mathbf{X}}\leftarrow{\mathbf{Y}}\rightarrow{\mathbf{A}}$, where graph structure and node features are both conditioned on node labels.
Definition 3.1.
(Feature Shift). Assume node features $x_{u},u\in{\mathcal{V}}$ are i.i.d sampled given labels $y_{u}$, then we have $\mathbb{P}({\mathbf{X}}|{\mathbf{Y}})=\prod_{u\in{\mathcal{V}}}\mathbb{P}(X_{u}|Y_{u})$. We then define the feature shift as $\mathbb{P}^{\mathcal{S}}(X_{u}|Y_{u})\neq\mathbb{P}^{\mathcal{T}}(X_{u}|Y_{u})$.
Definition 3.2.
(Structure Shift). As graph structure involves the connecting pattern between labels, we consider the joint distribution of the adjacency matrix and labels $\mathbb{P}({\mathbf{A}},{\mathbf{Y}})$, where Structure shift, denoted by $\mathbb{P}^{\mathcal{S}}({\mathbf{A}},{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}},{\mathbf{Y}})$, can be decomposed into as conditional structure shift (CSS) $\mathbb{P}^{\mathcal{S}}({\mathbf{A}}|{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}}|{\mathbf{Y}})$ and label shift (LS) $\mathbb{P}^{\mathcal{S}}({\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{Y}})$.
3.2 Theoretical Analysis
Let $\varepsilon^{\mathcal{S}}(g\circ\phi)$ denote the error of a pretrained GNN on the source domain and $\varepsilon^{\mathcal{T}}(g\circ\phi)$ the error of the model when applied to a test graph in the target domain. Inspired by Tachet des Combes et al. (2020), we provide an upper bound on the error gap between source and target domains, showing how a pretrained GNN (e.g., ERM) can be influenced. We denote the balanced error rate of a pretrained node predictor $\hat{Y}_{u}$ on the source domain as $\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\coloneq\max_{j\in\mathcal{Y}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}\neq Y_{u}|Y_{u}=j).$
Theorem 3.3 (Error Decomposition Theorem).
Suppose ${\mathcal{G}}^{{\mathcal{S}}}$ and ${\mathcal{G}}^{{\mathcal{T}}}$, we can decouple both graphs into independent ego-networks (center nodes and 1-hop neighbors). For any classifier $g$ with a mean pooling GNN encoder $\phi$ in node classification tasks, we have the following upper bound for on the error gap between source and target under feature shift and structure shift:
$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2\text{TV}(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}$
$+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Neighborhood shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature shift}}$
Our bound aligns with the findings of Liu et al. (2024), which highlight the impact of neighborhood shift, label shift, and feature shift on generalization to the target graph. We extend this understanding by deriving an explicit error bound. Notably, neighborhood shift is reduced from conditional structure shift given the assumptions in Thm.3.3.
4 Test-Time Structural Alignment (TSA)
@@
Motivated by our theoretical analysis, we propose TSA to address GTTA in a principled way. To address neighborhood shift, TSA first conducts neighborhood alignment via weighted aggregation and then properly balances self and neighboring node information combination based on the SNRs of these representations. Lastly, as a generic approach for neighborhood shift, TSA can be combined with non-graph TTA methods to refine the decision boundary to address the remaining feature and label shift.
@@ 4.1 Neighborhood Alignment
Neighborhood shift alters the node label ratios in the aggregated neighborhood information, causing the shift in the GNN-encoded representations. In training-time GDA, PairAlign (Liu et al., 2024) leverages such a technique by assigning edge weights to align the distribution of the source neighborhood with the target domain.
Inspired by this idea, for GTTA, we aim to achieve a similar goal but in a different direction. Based on Theorem 3.3, we align the target neighborhood distribution with the source domain, leveraging the fact that the pre-trained model is optimized for the source distribution. Specifically, the neighborhood distribution determines the ratio of messages passed from a neighboring class $j$ to center class $i$. To adjust for distributional differences, this ratio can be rescaled by assigning weights to edges from class $j$ to class $i$, effectively acting as an up/down-sampling mechanism during message aggregation. To ensure that message aggregation in the target domain aligns with the expected behavior in the source domain, TSA incorporates the following adjustment:
@@
To estimate $\bm{\gamma}$, we assume that the source summary statistics $\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ are recorded and available at test time; otherwise, alignment would not be feasible. Storing $\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ incurs minimal cost, as it is merely an $|{\mathcal{Y}}|\times|{\mathcal{Y}}|$ matrix. Beyond this, no additional information from the source domain is required. For $\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$, we estimate it based on target pseudo labels. Note that PairAlign (Liu et al., 2024) enhances estimation accuracy that relies pseudo labels (Liu et al., 2023) by leveraging a least-squares constrained optimization. However, in GTTA, the absence of source graphs and the potential need for real-time adaptation render this approach impractical.
While it seems direct to apply a similar alignment strategy by using the inverse of the ratio employed in PairAlign, assigning $\gamma$ requires knowledge of node labels, which is missing in the targe graph, making it challenging in GTTA.
@@
The ratio $\gamma$ should be assigned to edge weights based on the label pairs of the central and neighboring nodes. In training-time GDA, this assignment is straightforward as it relies on the ground-truth labels of the source graph. However, in GTTA, this information is unavailable. A naive approach is to use target pseudo labels, but this often results in significant mismatches. TSA addresses this by quantifying the uncertainty of target pseudo labels (Zhang et al., 2019b; Stadler et al., 2021; Hsu et al., 2022). In particular, TSA assigns $[\bm{\gamma}]_{i,j}$ only to node pairs $v\rightarrow u$ where both of their soft target pseudo labels $\hat{y}$ have low entropy $H(\hat{y})=-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{1}\cdot\ln(|{\mathcal{Y}}|)$. Here, $\rho_{1}$ is a hyperparameter and $\ln(|{\mathcal{Y}}|)$ is the maximum entropy in $|{\mathcal{Y}}|$ class prediction. In Fig.3 (a), nodes with low-entropy soft predictions are more reliable, resulting in higher accuracy after the assignment of $\bm{\gamma}$.
@@
4.2 SNR-Inspired Adjustment
Building on neighborhood alignment, we further optimize the signal-to-noise ratio (SNR) of node representations to enhance performance. Specifically, SNR is characterized by the ratio of inter-class representation distances to intra-class representation distances. A higher SNR indicates more informative and well-separated representations, which benefit classification.
Optimizing SNR complements the neighborhood alignment approach. Even if neighborhood label distributions are perfectly aligned, variations in neighbor set sizes between the source and target graphs can impact the SNR of aggregated neighboring node representations. Consequently, the combination of these aggregated representations with self-node representations in a typical GNN pipeline (Eq. 1) should be adjusted accordingly across source and target domains, particularly when the source and target graphs exhibit very different degree distributions. Furthermore, the SNR of node self-representations may vary across GNN layers, as deeper layers generally reduce variance. As a result, node self-representations in deeper layers tend to have higher SNR and should be assigned greater emphasis.
To implement the SNR-inspired adjustment, we introduce a parameter to perform the weighted combination of self-node representations and neighborhood-aggregated representations at each layer, adapting to node degrees as follows:
Definition 4.2.
Combined with $\bm{\gamma}$, $[\alpha]_{k}\cdot[\bm{\gamma}]_{i,j}$ is used to reweight the GNN message for non-self-loop node pairs, adjusting the influence from the node with a highly certain pseudo label $j$ to the node with a highly certain pseudo label $i$.
Remark. Neighborhood alignment alone does not address potential shifts in SNR. This is because the alignment approach, inspired by Thm. 3.3, focuses solely on expectation-based (first-order statistical) performance, whereas SNR also incorporates variance (second-order statistics). Thus, these two aspects complement each other.
@@ 4.3 Decision Boundary Refinement
In GTTA, label shift can result in a mismatch of the decision boundary, even after addressing neighborhood shift and obtaining high-SNR node representations. This is illustrated in Fig. 3 (b) and (c).
A straightforward approach to refining the decision boundary at test time is to adjust the classifier’s batch normalization parameters (TENT (Wang et al., 2021a)) or directly modify its output (T3A (Iwasawa & Matsuo, 2021) and LAME  (Boudiaf et al., 2022)). We integrate these techniques into our framework for two folds: (1) their refined pseudo-labels provide a more reliable assignment of $\bm{\gamma}$ and can supervise the update of SNR adjustment. (2) Reciprocally, better alignment of neighborhood information can further refine the decision boundary.
4.4 TSA Overview.
Note that $\bm{\gamma}$ is obtained from the initial pseudo labels. Only the parameters in $\text{MLP}^{(k)}$ and $b^{(k)}$ for $\alpha$ estimation in Eq. 3 need to be optimized in the test time according to Eq. 4.
@@
where $y_{u}^{\prime}$ is the hard pseudo label refined by procedure in Sec. 4.3 and $\hat{y}_{u}$ is the soft prediction from the original model. After updating $\alpha$, TSA makes predictions on the newly weighted graphs and then further adapts the boundaries as described in Sec. 4.3. Our proposed algorithm is summarized in Alg. 1.
Comparison to AdaRC (Bao et al., 2024) AdaRC is a recent work on GTTA that also integrates non-graph TTA methods with an approach for addressing structure shifts. However, it considers only degree shift and homophily shift, where homophily shift is merely a special case of neighborhood shift in our context. Consequently, AdaRC does not introduce the parameter $\bm{\gamma}$ and, therefore, lacks the ability to properly align neighborhood aggregated representations when the neighboring label distribution shifts.
@@ 5 Experiments
Table 4: Arxiv and DBLP/ACM no balanced training (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
@@
We evaluate TSA on synthetic datasets and 5 real-world datasets. More discussions and results in experiments can be found in Appendix D.
5.1 Datasets and Baselines
Synthetic Data. We use the CSBM model (Deshpande et al., 2018) to generate three-class datasets, focusing on structure shifts while keeping feature distributions unchanged. Specifically, we evaluate performance under three conditions: (1) neighborhood shift, (2) neighborhood shift plus SNR (induced by degree changes) shift, and (3) neighborhood shift combined with both SNR shift and label shift. This experimental setup is motivated by Thm. 3.3 and the observations in Sec.4.2. For each shift scenario, we examine two levels of severity, with the left column in later Table 2 corresponding to the smaller shift.
MAG (Liu et al., 2024) is a citation network extracted by OGB-MAG (Hu et al., 2020). Distribution shifts arise from partitioning papers into distinct graphs based on their countries of publication. The task is to classify the publication venue of each paper. Our model is pretrained on graphs from the US and China and subsequently adapted to graphs from other countries.
[…]
Arxiv (Hu et al., 2020) is a citation network between all Computer Science (CS) arXiv papers. Distribution shifts originate from different time. Our model is pretrained on the earlier time span 1950 to 2007/ 2009/ 2011 and test on later 2014 to 2016 and 2016 to 2018.
DBLP and ACM (Tang et al., 2008; Wu et al., 2020) are two citation networks. The model is trained on one network and adapted to the other to predict the research topic of a paper (node).
Baselines We compare TSA with six baselines. For non-graph TTA methods, we include TENT (Wang et al., 2021a), LAME (Boudiaf et al., 2022), and T3A (Iwasawa & Matsuo, 2021). Note that TENT is only applied to the classifier, as GNNs typically do not include batch normalization layers. GTrans (Jin et al., 2023), SOGA (Mao et al., 2024), and AdaRC (Bao et al., 2024) are direct comparisons in GTTA. AdaRC is limited to GPRGNN due to its design. We present the results for GraphSAGE (Hamilton et al., 2017) in the main paper, while the results for GPRGNN (Chien et al., 2020) (including AdaRC) are provided in the Appendix.
5.2 Result Analysis
From the MAG dataset results in Table 1, TSA combined with boundary refinement techniques consistently achieves top performance among all baselines. Compared to non-graph TTA baselines, such as TENT vs. TSA-TENT, TSA also provides considerable improvements over the original adaptation results. Notably, the benefits of TSA after neighborhood alignment are more pronounced in scenarios with larger neighborhood shifts, such as adaptation from the US to FR, as reflected in the dataset shift metrics (Table LABEL:table:magstats).
Among all baselines, T3A consistently outperforms as a representative non-gradient-based method that refines the decision boundary, while TENT struggles in most scenarios. This is primarily due to the relatively imbalanced class distribution in the MAG datasets, where evaluation is conducted on the top 19 classes. Since TENT learns feature-wise transformations through entropy optimization on the target graph, it tends to be biased toward the dominant class. Other GTTA methods, such as GTrans, achieve performance comparable to ERM, whereas SOGA sometimes underperforms due to its strong reliance on the graph homophily assumption, which is often weak in sparse graphs with many classes, such as MAG and Arxiv.
The results for the synthetic datasets are presented in Table 2. The first six columns correspond to experiments conducted under imbalanced source training. Compared to real datasets, we observe more significant advantages from TSA. TENT outperforms LAME and T3A, as it directly maximizes the most probable class through entropy minimization. In contrast, LAME incorporates regularization, while T3A relies on distance to prototypes, making it less sensitive to majority-class dominance (Zhang et al., 2022). When combined with TSA, we observe similar results across all three variants, indicating that the primary adaptation ability stems from TSA rather than non-graph TTA methods. While GTTA baselines perform better than ERM, they still fall short of our TSA-based methods. TSA also demonstrates strong performance under additional shifts in SNR caused by degree shifts. Additionally, we consider training a pretrained model on a balanced source domain under the same edge connection probability as the previous two columns. The last two columns demonstrate better performance than the previous two under balanced training, even though they encounter the same shift in label connection. This showcases that dataset imbalance is a severe issue in source training, which aligns with the worst-case error described in Theorem 3.3.
The pileup results in Table 3 demonstrate the effectiveness of TSA-based methods under both neighborhood shift and label shift, with each module contributing in different ways. The neighborhood alignment module provides significant improvements when generalizing from a low PU level to a high PU level, as this scenario is primarily dominated by neighborhood shift. Conversely, when adapting from a high PU level to a low PU level, label shift becomes the dominant factor. In this case, incorporating non-graph TTA techniques is crucial, while neighborhood alignment offers an additional but smaller benefit. Moreover, TENT can be vulnerable when generalizing from an imbalanced source graph to a balanced target graph. This is because source training on high PU graphs suffers from class imbalance, making entropy minimization non-robust due to a mismatched decision boundary when adapting to low PU graphs. The last two columns, which represent the same PU level but different physical signals, show minimal benefits from TSA. This is expected, as these scenarios primarily exhibit feature shifts rather than structure or label shifts.
The results from Arxiv and DBLP/ACM in Table 4 confirm that TSA can be effectively integrated with non-graph TTA methods to mitigate feature shift while also facilitating reasonable adjustments for structure shifts. These citation networks exhibit relatively mild structure shifts, leading to closely comparable performance across all baselines. TENT and TSA-TENT stand out by maximizing the majority classes through entropy minimization. Nonetheless, we still observe meaningful improvements with the addition of TSA when compared to the corresponding baselines, demonstrating its effectiveness in enhancing adaptation.
5.3 Analysis of $\alpha$ for SNR-induced refinement
@@
As defined in Eq.3, $\alpha$ is learned, depending on node degrees and GNN layer depth. The trend of $\alpha$ in Fig. 3 (d) aligns with our expectations: earlier GNN layers require less attention to self-node representations (small $\alpha$), while deeper layers increasingly rely on them. Additionally, the plot reveals a correlation with node degrees, where nodes with higher degrees (and thus higher SNR) place greater weight on neighborhood-aggregated messages. However, the overall effect of different GNN layers appears to be greater than that of varying node degrees.
5.4 Ablation Studies
In Table 6 (Appendix), we examine the effects of neighborhood alignment and SNR-based representation adjustments. The results consistently show that incorporating both modules yields the best performance. In some cases, removing SNR adjustments leads to a greater performance drop compared to removing neighborhood alignment. However, this does not necessarily indicate that neighborhood alignment is less important. Instead, we think this effect arises because the additional backpropagation training for SNR adjustments amplifies their contributions.
6 Discussion and Limitations

### v2

@@ 4 Test-Time Structural Alignment (TSA)
Figure 3: (a) Comparison of neighborhood alignment with $\bm{\gamma}$ from model prediction and Oracle on the CSBM graphs [60]. The top (or bottom) subfigures represent the assignment under neighbor shift (or neighbor shift plus label shift, respectively). Nodes are grouped by the entropy of their soft pseudo labels and the y axis shows the accuracy after assigning $\bm{\gamma}$. Ideally, a correct assignment (red) would lead to near-perfect accuracy. However, the assignment based on pseudo labels is far from optimal (blue). Figure (b) to (d) present t-SNE visualizations of node representations during GTTA. The model is trained on the source domain in (b) (CSBM with label distribution $[0.1,0.3,0.6]$) and evaluated on the target domain in (c) (CSBM with label distribution $[0.3,0.3,0.3]$). Figure (d) illustrates the qualitative result of the neighborhood alignment. Despite well-aligned neighborhood representations, a mismatched decision boundary (red) leads to class 0 misclassifications. Node colors represent the ground-truth labels. The top subfigures show the output given by the GNN encoder, while the bottom subfigures show the classifier decision boundaries.
@@
Motivated by our theoretical analysis, we propose TSA to address GTTA in a principled way. To address structure shift, TSA first conducts neighborhood alignment via weighted aggregation and then properly balances self and neighboring node information combination based on the SNRs of these representations. Lastly, as a generic approach for structure shift, TSA can be combined with non-graph TTA methods to refine the decision boundary.
@@ 4.1 Neighborhood Alignment
The conditional structure shift (CSS) alters the node label ratios in the aggregated neighborhood information, causing a shift in the GNN-encoded representations. Based on Theorem 3.3, we align the target neighborhood distribution with the source domain, leveraging the fact that the pre-trained model is optimized for the source distribution. Specifically, the neighborhood distribution determines the ratio of messages passed from a neighboring class $j$ to the center class $i$. To adjust for distributional differences, this ratio can be rescaled by assigning weights to edges from class $j$ to class $i$, effectively acting as an up/down-sampling mechanism during message aggregation. To ensure that message aggregation in the target domain aligns with the expected behavior in the source domain, TSA incorporates the following adjustment:
@@
To estimate $\bm{\gamma}$, we estimate $\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ based on target pseudo labels. $\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ can be computed directly from the source graph during model training and added as extra model parameters for the test time direct usage. During test time, we weight the message with a properly assigned $[\bm{\gamma}]_{i,j}$ computed based on the ratio of the two quantities. This idea aligns with the general principle of test-time alignment [43, 45, 46], but is novel in the aspect of dealing with the structure shift for graph data. Assigning edge weights to reduce structure shifts was also studied in training-time GDA by PairAlign [59]. TSA differs from PairAlign in that $\bm{\gamma}$ estimation and model training can be done together in PairAlign, which better coordinates the two sides as it does training-time GDA (the model better fits the weighted message passing, and the weights are adjusted based on the improved models). However, this cannot be done at test time. In particular, to compensate for such a loss, TSA needs a more reliable assignment strategy specified in the following.
@@
The ratio $\bm{\gamma}$ should be assigned to edge weights based on the label pairs of the central and neighboring nodes. In training-time GDA, this assignment is straightforward, as it relies on the ground-truth labels of the source graph. However, in GTTA, the node labeling information is unavailable. A naive approach is to use target pseudo labels, i.e., $[\bm{\gamma}]_{i,j}$ for the message from a node with pseudo label $j$ to a node with pseudo label $i$, but this often results in significant mismatches. TSA addresses this by quantifying the uncertainty of target pseudo labels [61, 62, 63, 64]. In particular, TSA assigns $[\bm{\gamma}]_{i,j}$ only to node pairs $v\rightarrow u$ where both of their soft target pseudo labels $\hat{y}$ have low entropy $H(\hat{y})=-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{1}\cdot\ln(|{\mathcal{Y}}|)$. Here, $\rho_{1}$ is a hyperparameter and $\ln(|{\mathcal{Y}}|)$ is the maximum entropy for a $|{\mathcal{Y}}|$ class prediction. In Fig. 3 (a), nodes with low-entropy soft predictions are more reliable, resulting in higher accuracy after the assignment of $\bm{\gamma}$.
[…]
Refer to caption Refer to caption Refer to caption
Figure 4: (a) and (b) compare the SNR of neighborhood-aggregated representations to that of self-node representations across node degrees and GNN layers, based on the ratio of the two quantities. A higher ratio indicates that the neighborhood-aggregated representations exhibit higher SNR. (c) Analysis of the SNR adjustment $\alpha$ across different layers and node degrees. Details are discussed in Sec. 5.2.
4.2 SNR-Inspired Adjustment
Building on neighborhood alignment, we further optimize the signal-to-noise ratio (SNR) of node representations to enhance performance. Specifically, the SNR is characterized by the ratio of inter-class representation variance to intra-class representation variance.
$\text{SNR}\coloneqq\frac{\sigma_{inter}^{2}}{\sigma_{intra}^{2}},\quad\sigma_{inter}^{2}=\frac{1}{|{\mathcal{Y}}|}\sum_{i\in{\mathcal{Y}}}n_{i}\left\|\mu_{i}-\mu_{*}\right\|_{2}^{2},\quad\sigma_{intra}^{2}=\frac{1}{|{\mathcal{Y}}|}\sum_{i\in{\mathcal{Y}}}\sum_{h\in\mathcal{H}_{i}}\left\|h-\mu_{i}\right\|_{2}^{2}$ (3)
where $n_{i}$ is the number of samples in class $i$, $\mu_{i}$ is the centroid representation of class $i$, $\mu_{*}$ is the overall centroid of all nodes, and $\mathcal{H}_{i}$ represents the set of representations in class $i$. A higher inter-class variance $\sigma_{inter}^{2}$ indicates informative and well-separated representations, while a higher $\sigma_{intra}^{2}$ indicates a higher noise in the representations.
Optimizing SNR complements the neighborhood alignment approach. Even if neighborhood label distributions are perfectly aligned, variations in neighbor set sizes between the source and target graphs can impact the SNR of aggregated neighboring node representations. To compare the SNR of neighborhood-aggregated representations to that of self-node representations, we compute SNR ratio=Nbr. Aggr. SNRSelf-Node $\text{SNR ratio}=\frac{\text{Nbr. Aggr. SNR}}{\text{Self-Node SNR}}$ which exhibits different trends as a function of node degree (aka neighbor set size) and layer depth, as shown in Fig. 4 (a) and (b). In both figures, deeper layers and higher degree values generally lead to a higher SNR ratio, indicating aggregated representations are better denoised. However, the rate and shape of SNR ratio growth vary across source and target domains. Consequently, the combination of these aggregated representations with self-node representations in a typical GNN pipeline (Eq. 1) should be properly adjusted as well in test-time adaptation. To implement the SNR-inspired adjustment, we introduce a parameter to perform the weighted combination of self-node representations and neighborhood-aggregated representations at each layer, adapting to node degrees as follows:
Definition 4.2.
Together, $[\alpha]_{k}\cdot[\bm{\gamma}]_{i,j}$ is used to reweight the GNN messages, adjusting the influence from the node with a highly certain pseudo label $j$ to the node with a highly certain pseudo label $i$.
Remark. Neighborhood alignment alone does not address potential shifts in SNR. This is because the alignment approach, inspired by Thm. 3.3, focuses solely on population-level performance, whereas SNR also incorporates the noise (variance) induced by empirical sampling from the population. Thus, these two aspects complement each other.
@@ 4.3 Decision Boundary Refinement
In GTTA, even when CSS is fully addressed and high-SNR node representations are obtained, the decision boundary may still suffer from a mismatch due to label and feature shifts (Fig. 3 (d)).
A straightforward approach to refining the decision boundary at test time is to adjust the classifier’s batch normalization parameters (TENT [28]) or directly modify its output (T3A [26] and LAME  [27]). We integrate these techniques into our framework for two folds: (1) their refined pseudo-labels provide a more reliable assignment of $\bm{\gamma}$ and can supervise the update of SNR adjustment. (2) Reciprocally, better alignment of neighborhood information can further refine the decision boundary.
Matcha [65] is a very recent work on GTTA that also integrates non-graph TTA methods with an approach for addressing structure shift. However, it considers only degree shift and homophily shift, where homophily shift is merely a special case of CSS in our context. Consequently, Matcha does not introduce the parameter $\bm{\gamma}$ and does not fully address the structure shift.
TSA Overview. Note that $\bm{\gamma}$ is obtained from the initial pseudo labels. Only the parameters in $\text{MLP}^{(k)}$ and $b^{(k)}$ for $\alpha$ estimation in Eq. 4 need to be optimized at test time according to Eq. 5.
@@
where $y_{u}^{\prime}$ is the hard pseudo label refined by the procedure in Sec. 4.3 and $\hat{y}_{u}$ is the soft prediction from the original model. After updating $\alpha$, TSA makes predictions on the newly weighted graphs and then further adapts the boundaries as described in Sec. 4.3. Our algorithm is summarized in Alg. 1.
@@ 5 Experiments
We evaluate TSA on synthetic datasets and 5 real-world datasets. Additional results and discussions (e.g., hyperparameter sensitivity) are provided in Appendix D.
5.1 Datasets and Baselines
Synthetic Data. We use the CSBM model [60] to generate three-class datasets, focusing on structure shifts while keeping feature distributions unchanged. Specifically, we evaluate performance under three conditions: (1) CSS, (2) CSS plus SNR (induced by degree changes) shift, and (3) full structure shift (CSS + label shift) plus SNR shift. This experimental setup is motivated by Thm. 3.3 and the observations in Sec.4.2. For each shift scenario, we examine two levels of severity, with the left column in Table 1 corresponding to the smaller shift.
MAG [59] is a citation network extracted by OGB-MAG [66]. Distribution shifts arise from partitioning papers into distinct graphs based on their countries of publication. The task is to classify the publication venue of each paper. Our model is pretrained on graphs from the US and China and subsequently adapted to graphs from other countries.
@@ 5.1 Datasets and Baselines
Arxiv [66] is a citation network between all Computer Science (CS) arXiv papers. Distribution shifts originate from different time periods. Our model is pretrained on the earlier time span 1950 to 2007/ 2009/ 2011 and tested on later periods 2014-2016 and 2016-2018.
DBLP and ACM [68, 5] are two citation networks. The model is trained on one network and adapted to the other to predict the research topic of a paper (node).
Baselines. We compare TSA with six baselines. For non-graph TTA methods, we include TENT [28], LAME [27], T3A [26], and ActMAD [43]. Note that TENT is applied only to the classifier, as GNNs typically do not include batch normalization layers. GTrans [30], HomoTTT [31], SOGA [32], and Matcha [65] are direct comparisons in GTTA. Matcha is limited to GPRGNN due to its design. We present the results for GraphSAGE [22] in the main paper, while the results for GPRGNN [69] (including Matcha) and GCN [20] are provided in the Appendix.
5.2 Result Analysis
@@
Table 4: Arxiv and DBLP/ACM results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
@@
Comparison with Baselines. TSA achieves strong performance across all datasets and GNN backbones. In all cases, the best results are obtained by GTTA methods (underlined across Tables 14), highlighting the importance of addressing structure shift in graph data. In the synthetic dataset (Table 1), TSA outperforms GTTA baselines under plain CSS, demonstrating neighborhood alignment as an effective approach to handling CSS. The same table also showcases that dataset imbalance severely affects source training, which aligns with the worst-case error described in Theorem 3.3. In real-world datasets, TSA achieves an average improvement of 10% over existing GTTA baselines. GTrans relies on self-supervised learning to augment graphs. Although it shows improvements on CSBM, its performance is limited on real-world datasets (Tables 24). SOGA and HomoTTT depend on the homophily assumption, which does not always hold in real-world graphs [70], leading to unstable performance. In contrast, TSA makes no such assumptions and consistently delivers stable improvements by structure shifts.
Different TSA-version. Overall, TSA-variants consistently yield performance gains of up to 21.25% over the corresponding non-graph TTA baselines (bold across Table 1- 4). In particular, the performance of the TSA variants differs in how the boundary refinement addresses the label shift. TSA-T3A is less sensitive to the dominance of the majority class, as it refines the boundary based on distances to prototypes, offering increased robustness [71]. Therefore, TSA-T3A works well on imbalanced datasets such as MAG (Table 2) and Pileup (Table 3). We observe TSA-TENT and TSA-LAME perform well on Arxiv and ACM/DBLP (Table 4), as they directly maximize the most probable class, with TSA-LAME introducing a regularization overhead.
@@
Analysis of $\alpha$. The learned $\alpha$ depends on node degrees and GNN layer depth. The trend of $\alpha$ in Fig. 4 (c) aligns with our expectations: earlier GNN layers require less attention to neighborhood-aggregated representations (small $\alpha$), while deeper layers increasingly rely on them. Additionally, the plot reveals a correlation with node degrees, where nodes with higher degrees (and thus higher SNR ratios) place greater weight on neighborhood-aggregated messages. However, the overall effect of different GNN layers appears to be greater than that of varying node degrees. Lastly, the learned $\alpha$ exhibits a trend similar to the source SNR ratio shown in Fig. 4 (a), assigning lower weights in layer 1 and nearly identical weighting patterns in layer 2 and 3.
Ablation Studies. In Table 5 (Appendix), we examine the gains of neighborhood alignment and SNR-based representation adjustments on both synthetic and real-world datasets. The results consistently show that incorporating both modules yields the best performance. Removing neighborhood alignment leads to a performance drop of up to 19.9% on the synthetic dataset and 3.16% on the real-world dataset, while removing SNR adjustments results in a drop of up to 2.43% and 2.97%, respectively. Removing neighborhood alignment causes a larger drop in CSBM due to the more significant shift used in the data modeling.
Computational Efficiency. In Table 6 (Appendix), we quantify the computation time of TSA-T3A on MAG US, containing the largest number of nodes and edges. The time complexity of neighborhood alignment is $O(1)$ as it is a one-step edge weight assignment. SNR adjustment contributes relatively more to the overhead due to backpropagation. Notably, compared to GTTA baselines, TSA-T3A is approximately 2, 4 and 8 times more efficient than HomoTTT, GTrans and SOGA.
6 Conclusion